# LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues

### v1

@@ 4.1 Diversity of Slots and Intents
The generated LUCID data contains more intents and slots than existing task-oriented dialogue datasets (Table 1). Specifically, the dataset contains 100 intents, across 13 domains, with 501 different slots. While the SGD dataset contains more domains than LUCID, these domains are narrower in scope. For example, SGD includes separate domains for buses, taxis, flights and trains, while LUCID has a single transportation domain incorporating intents for each of these areas. The additional number of slots and intents in LUCID illustrates our ability to create diverse and challenging data using LUCID, despite generating a smaller dataset compared to SGD and MultiWOZ (see Table 4). Considering the size of our dataset, compared to SGD and MultiWoz there are almost twice as many unique slot values, demonstrating the level of variety in our dataset (see values per turn in Table 4).
We compare the number of slots within LUCID to PRESTO, SGD and MultiWoZ in Table 1, showing that after excluding duplicated slots in both PRESTO and SGD, the LUCID generated data has almost twice the number of slots. Appendix F outlines our approach for removing these duplicated slots, in the case of SGD when the same slots are included in different services, or in PRESTO when slots are counted multiple times if they have slightly different parse trees.
@@
Table 1: Summary statistics of our dataset, displaying the number of domains and intents present, the number of intents per domain, the number of slots present, and the number of explicitly labelled conversational phenomena (unhappy paths). $\dagger$ considered the 303 slots in the English intents. Unlike Table 4, this table considers all splits in the dataset.
It is worth pointing out that the LUCID dataset was generated primarily to showcase the capabilities of the LUCID data generation system. Others are free to use the LUCID system to generate much larger and even more complex datasets, with many more domains, intents and interesting conversational phenomena. This extensibility is what most clearly distinguishes LUCID from these other data generation efforts.
@@ 4.2 Conversational Phenomena
LUCID contains a greater number of labeled conversational phenomena than existing dialogue datasets (Table 1). The recently released PRESTO dataset also contains turn-level annotated phenomena, labelling six types of unhappy paths (Goel et al., 2023). These unhappy paths include in-turn corrections, correcting actions, correcting slot values, code-mixing, disfluencies and cancellations. While half of these phenomena relate to corrections, only two of our labeled phenomena are for corrections, correcting slot values either in-turn or across turns. Instead, we focus on distinguishing between relevant, sensible user replies and where a virtual assistant should ask for clarification rather than using the initial response to populate slots.
While not explicitly stated, some specific conversational phenomena can also be derived from existing datasets, such as co-references Cheng et al. (2020), or user corrections Rastogi et al. (2020).
@@ 4.3 Qualitative and Quantitative Analysis
We perform a qualitative analysis on the generated dataset, identifying issues that can be addressed by filtering out more examples in a post-processing step. Informed by this analysis, we perform a further quantitative analysis of 200 conversations in our dev set, measuring the quality of the system labels and investigating the impact of any identified issues. See Appendix A for the full results of our quantitative analysis, and Appendix B for details of additional post-processing. By publishing these more detailed quality metrics, we aim to raise the bar for future data generation efforts.
As a related issue, the natural language response can also suggest that an intent has been performed before confirmation is given by the user. Informed by this finding, we filter out conversations when an intent was not performed, unless there was a cancellation signal provided by the user (removing 79 conversations). After this filtering, our quantitative analysis finds that 4.5% of conversations include responses that suggest an intent has been performed before confirmation is given. However, in each case there was no impact on the conversation beyond the phrasing of the natural language response. As LUCID prioritises validating the system labels, there is room for improvement in future work to also improve the corresponding natural language responses.
Finding 2) The planner’s choice of slots and their corresponding values can sometimes be unrealistic. While a strength of LUCID is the realistic and varied slot values used in conversations, this is not always the case. We also notice that the choice of slots included in a conversation is not always realistic. For example, you would not usually give the start time, end time in addition to specifying the duration of a swimming lesson. Our quantitative analysis identifies that 6% of conversations contain at least one unrealistic slot value, while 3.5% of conversations include an unrealistic combination of slots. The unrealistic slot combinations demonstrate a limitation to our sampling approach, where we randomly sample which optional slots should be included in each conversation.
Finding 4) The system command labelling is consistently high quality, with few labelling mistakes. In a quantitative analysis of 200 conversations in the development set, we find only 1% of examples when the system label is not correct. In comparison, Eric et al. (2020) identify annotation errors in 40% of turns in MultiWoz 2.0, demonstrating the relative quality of the LUCID system labels. More detail on the two system labelling errors identified are provided in Appendix A.
Additionally, we find that only 8.5% of conversations have a string slot value that does not match the reviewer’s first choice. However, these differences are subjective and subtle, with over half of the cases concerning whether to include ‘the’ before a date (e.g. ‘the 5th of March’ vs ‘5th of March’).
@@ 5 Baseline Results
We train T5 Raffel et al. (2020) and Flan-T5 Chung et al. (2022) baseline models on LUCID, evaluating on both our in-distribution and out-of-distribution test sets. When retrieving intent schemas, a Sentence-BERT Reimers and Gurevych (2019) model is used to encode the tool name and the last user utterance, choosing the tool with the highest cosine similarity. As gold system labels for previous turns are seen in the prompt conversation history, we retrieve all tools that have been mentioned in an oracle history up to that point.
As expected, the joint goal accuracy is considerably higher when evaluating on the seen test set compared to the unseen test set, with accuracy scores of 67.5% and 42.2% respectively for a T5-base model (see Table 5). We also isolate the impact of the retrieval model, comparing three scenarios: 1) using our tool retrieval, 2) using an oracle tool retrieval, and 3) including no tools in the prompt. We find that the tool retrieval is not a major weakness of our baseline model (see Table 6).
Finally, we evaluate our Flan-T5-base model on each different conversational phenomena (see Section 3.5), highlighting sarcasm, ASR-early end examples, and answering about a different slot as the most challenging phenomena. On the other hand, cancellations, users delaying giving confirmation, and overheard answers are more easily predicted, and our baseline model makes no mistakes in these cases. Full details of our evaluation metrics, training setup and prompts can be found in Appendix C.
@@ 6 Conclusion
We introduce LUCID, a pipeline of LLM calls which is designed to create high quality and linguistically sophisticated dialogue data. LUCID involves an extensive validation process, including three validator LLMs that discard conversations where there is any disagreement. To demonstrate the quality of the data produced, we generate a seed dataset of 4,277 dialogues, consisting of 92,699 turns, with a wide variety of challenging conversational phenomena. These phenomena are all labeled at a turn level. The generated system labels in LUCID prove to be highly accurate, with only 1% of conversations containing system errors. This allows for reliable evaluation across our nine different conversational phenomena, either on seen or unseen intents. We also provide a detailed analysis of our generated data, identifying areas that could be improved in future data generation efforts. Finally, we make our code available to facilitate larger scale, high quality data generation with more intents, domains and complex conversational phenomena.
Appendix A Quantitative Analysis System Label Noise
Table 2 provides full results from our quantitative analysis of 200 dev examples. In this section, we additionally explain the issues that caused the two incorrect system labels identified.
@@ Appendix A Quantitative Analysis System Label Noise
Appendix B Additional Post Processing
We provide more detail on the other filters that were used in post processing to improve the quality of the data.
We filter conversations where a slot being is corrected during the conversation, but there is no correctional signal provided by the user, used for our complex phenomena validation (mentioned in Section 3.5). This filters out instances where a slot was first mentioned by the user without giving a value, with the system incorrectly assigning a slot value from this turn. The user will then follow-up in another turn with the desired slot value. This new validation filters out 123 conversations, which entirely resolves this issue.
We perform additional filtering to remove empty string slot values (removing 27 conversations), and any instances where the system turn predicts a hint, as hints should only occur in Signal turns. There were 172 occurrences when a hint was predicted by the system, although in almost all cases these conversations were already filtered by another post-processing filter.
@@ Appendix B Additional Post Processing
Appendix C Baseline results
@@
We train six different baseline models on the LUCID training data (T5-small, T5-base, T5-large, Flan-T5-small, Flan-T5-base and Flan-T5-large models). Each model is evaluated on the test set for seen intents and the OOD test set for unseen intents (see Table 5). We additionally experiment with training our Flan-T5-base baseline on varying amounts of training data (see Table 7). Details about about the choice of hyper-parameters can be found in Appendix I.
@@
To measure performance on our generated LUCID data, we consider five performance metrics: joint goal accuracy, intent accuracy, fuzzy slot accuracy, exact match accuracy between user turns, and exact match accuracy for an entire dialogue.
@@
We also use fuzzy matching for our slot accuracy measure, which is a joint accuracy measure across all the slot values provided in a single system turn (when any slot values are predicted, or when they are included in the labels). We additionally introduce exact match metrics that consider the accuracy of all system commands, not just those that refer to intent and slot values (for example, including ‘say’ commands). We introduce two exact match metrics - exact match (turn) considers whether all predicted system commands between two user turns exactly match their labels, while exact match (conversation) considers whether every predicted system command in a conversation matches with the system labels. We introduce the exact match metrics as a result of the consistently high quality string slot values in the generated data (see Section 4.3).
@@
Table 3: Total number of each conversational phenomenon within each split of our dataset. While there are few conversations for cancellation, this behaves similarly to the ‘delay confirmation’ phenomenon. We also show the exact match (Turn) metric for each conversational phenomena from a T5-Flan-base baseline model (details of the metric are provided in section Appendix C.
Appendix D Intent generation
@@
Additionally, query intents were merged together for intents that had the same generated entity names. As a result, there are 6 fewer query intents than transactional intents, resulting in a total of 100 intents. Each of the following pairs of transactional intents returned the same entity names, and so their corresponding query intents were combined: add_tv_program_to_favorites and add_artist_to_favorites (both of which return ‘favorites’ entities), set_timer and set_alarm (both of which return ‘alarm’ entities), book_nails_appointment and book_spa_appointment (both of which return ‘appointments’), order_supermarket_shop and order_takeaway (both of which return ‘orders’), add_song_to_favorites and play_song (both of which return ‘songs’), and review_film and review_restaurant (both of which return ‘reviews’).
Appendix E Slot Value Generation
This section provides more detail to describe how the conversation planner creates slot values, using a range of different stages.
For every additional intent in the conversation, an LLM generates a plausible reason why the user would want to complete this intent given what has already occurred in the conversation (stage 5). Likely slot values are then generated (stage 6) based on this context. Stage 4 is then repeated, increasingly the likelihood that the slot values selected are realistic and coherent for the intent. An LLM is then asked to update the slot values across every intent in the conversation so that these intents are related and consistent (stage 7), encouraging more natural conversations that align more closely with human to virtual assistant interactions.
@@ Appendix E Slot Value Generation
Appendix F Slot duplication within PRESTO and SGD

### v2

@@ 4.1 Diversity of Slots and Intents
The generated LUCID data contains more intents and slots than existing task-oriented dialogue datasets (Table 1). Specifically, the dataset contains 100 intents, across 13 domains, with 501 different slots. While the SGD dataset contains more domains than LUCID, these domains are narrower in scope. For example, SGD includes separate domains for buses, taxis, flights and trains, while LUCID has a single transportation domain incorporating intents for each of these areas. The larger number of slots and intents in LUCID illustrates our ability to create diverse and challenging data using LUCID, despite generating a smaller dataset compared to SGD and MultiWOZ (see Table 5).
@@
Table 1: Summary statistics of our dataset, displaying the number of domains and intents present, the number of intents per domain, the number of slots present, and the number of explicitly labelled conversational phenomena (unhappy paths). For PRESTO, we consider the 303 slots in English intents ($\dagger$). Unlike Table 5, this table considers all splits in the dataset. Appendix I describes how duplicate slots and intents are removed for SGD and PRESTO.
As the LUCID dataset was generated primarily to showcase the capabilities of the LUCID data generation system, others are free to use the LUCID system to generate much larger and even more complex datasets. This extensibility is what most clearly distinguishes LUCID from these other data generation efforts.
@@ 4.2 Conversational Phenomena
LUCID contains a greater number of labelled conversational phenomena than existing dialogue datasets (Table 1). The recently released PRESTO dataset also contains turn-level annotated phenomena, labelling six types of unhappy paths (Goel et al., 2023). These unhappy paths include in-turn corrections, correcting actions, correcting slot values, code-mixing, disfluencies and cancellations. While half of these phenomena relate to corrections, this is the case for only two of our labelled phenomena, correcting slot values either in-turn or across turns. Instead, we focus on distinguishing between relevant, sensible user replies from cases where a virtual assistant should ask for clarification (rather than using the initial response to populate slot values).
@@ 4.3 Qualitative and Quantitative Analysis
We perform a qualitative analysis on the generated dataset (conducted by one of the paper authors) to thoroughly investigate the dataset quality and identify any issues. This included a manual review of 200 conversations in our dev set, which only highlighted two labelling errors (impacting only 1% of conversations). In comparison, Eric et al. (2020) identify annotation errors in 40% of turns in MultiWoz 2.0, demonstrating the relative quality of the LUCID system labels.
The two labelling errors identified in this review involved: 1) A user mentioning there will be no spoilers in a review, where LUCID correctly assigns the spoiler alert slot value as False, but additionally includes the text ‘no spoilers in my review’ as part of the review itself. 2) LUCID not recognising an in-turn correction by the user, mistakenly including all of the user text (including the correction itself) as part of a slot value. See Appendix C for further details. We additionally share a qualitative analysis of our data in Appendix D, highlighting specific areas where our method could be further improved. We include this analysis to further raise the bar for future LLM data generation efforts.
@@ 5 Baseline Results
Intent acc. JGA
Test (seen):
T5-Small 94.7 57.1
T5-Base 97.9 67.5
T5-Large 98.7 69.0
Test-OOD (unseen):
T5-Small 95.3 22.0
T5-Base 97.6 42.2
T5-Large 98.8 45.7
Table 2: Results of our baseline model (with retrieval). Full results and evaluation metric descriptions are provided in Appendix G.
We train T5 Raffel et al. (2020) and Flan-T5 Chung et al. (2022) baseline models on LUCID, evaluating on both our in-distribution and out-of-distribution test sets. When retrieving intent schemas, a Sentence-BERT Reimers and Gurevych (2019) model is used to encode the tool name and the last user utterance, choosing the tool with the highest cosine similarity.
As expected, the joint goal accuracy is considerably higher when evaluating on the seen test set compared to the unseen test set, with accuracy scores of 67.5% and 42.2% respectively for a T5-base model (see Table 6). We also isolate the impact of the retrieval model, comparing three scenarios: 1) using our tool retrieval, 2) using an oracle tool retrieval, and 3) including no tools in the prompt. We find that the tool retrieval is not a major weakness of our baseline model (see Table 7). Finally, we evaluate our Flan-T5-base model on each different conversational phenomena (see Section 3.5), highlighting sarcasm, ASR-early end examples, and answering about a different slot as the most challenging phenomena. Full experimentation details and results can be found in Appendix G.
@@ 6 Conclusion
We introduce LUCID, a pipeline of LLM calls which is designed to create high quality and linguistically sophisticated dialogue data. LUCID involves an extensive validation process, including three validator LLMs that discard conversations where there is any disagreement. To demonstrate the quality of the data produced, we generate a seed dataset of 4,277 dialogues, consisting of 92,699 turns, with a wide variety of challenging conversational phenomena. The generated system labels in LUCID prove to be highly accurate, with only 1% of conversations containing a labelling error. We make our code available to facilitate larger scale, high quality data generation.
Limitations
The main limitation of our approach is the cost of using a closed-source LLM. This prevented us from generating a larger number of dialogues or performing more ablation studies to isolate the improvements from specific stages. This cost was driven by our if in doubt, discard approach to validation, which prioritised the accuracy and quality of the data produced, at the expense of the computational time and cost involved. While there are also substantial costs associated with high quality manual annotation, in this work we aim to show that an LLM-driven approach to generating high quality data is possible and feasible. We also aim to produce a seed dataset of the highest quality which can be used by practitioners on an on-going basis.
Appendix A Related work - Data Generation with LLMs
Wu et al. (2023) recently introduce a framework allowing the interaction of multiple, different LLMs, based on the idea that LLMs can solve highly challenging tasks if these tasks are broken into smaller steps. While Wu et al. (2023) are successful in generating dialogues for a group chat scenario, this does not require the intent and slot labelling needed for task-oriented dialogue. For other NLP tasks, to avoid labelling errors or poor quality data, generating data with LLMs can involve human annotators reviewing the generated utterances Liu et al. (2022); Wiegreffe et al. (2022), or using the generated data as unlabelled data to be used with knowledge distillation Meng et al. (2022); He et al. (2022); Stacey and Rei (2023). Labelled data generation has been successful for other tasks without human input Honovich et al. (2023); Ye et al. (2022); Schick and Schütze (2021); West et al. (2022); Wu et al. (2022), however noise may be an issue for a large proportion of the data Honovich et al. (2023); Schick and Schütze (2021).
Appendix B Conversation Planner Details
To ensure variety in the generated conversations, the planner makes extensive use of sampling, choosing how many intents should be provided in the conversation, which optional slots should be discussed, the conversational phenomena (both happy and unhappy paths) that should be included, and which slots and intents any conversational phenomena should be applied to.
Sampling is also involved in choosing the path of intents that are included in the conversation. The planner is provided with a sample of intents across all domains, before an LLM generates a plausible sequence of intents from this sample (stage 3).
We also provide more detail below to describe exactly how the conversation planner creates slot values for the conversation, involving a range of different stages:
For all other intents in the conversation (after the first intent), an LLM generates a plausible reason why the user would want to complete this intent given what has already occurred in the conversation (stage 5). Likely slot values are then generated (stage 6) based on this context. Stage 4 is then repeated, increasingly the likelihood that the slot values selected are realistic and coherent for the intent. An LLM is then asked to update the slot values across every intent in the conversation so that these intents are related and consistent (stage 7), encouraging more natural conversations that align more closely with human to virtual assistant interactions.
Appendix C Quantitative Analysis Findings
Table 3 provides the full results from our quantitative analysis of 200 dev examples. This analysis identified two labelling issues within the 200 conversations that were reviewed (conducted by one of the paper authors).
@@ Appendix C Quantitative Analysis Findings
Table 3 also measures some of the issues reported in the qualitative analysis, including: 1) how many times the natural language generated response (NLG) indicates an intent was performed before confirmation was given, 2) when the NLG does not follow the system predictions, 3) unrealistic slot values, 4) unrealistic combination of slots mentioned for an intent, 5) when the first choice string span was not selected by LUCID, and 6) when the user or NLG does not fully understand the purpose of the intent. While previous work does not report similar metrics, we publish these figures with the aim of raising the bar for future data generation efforts.
In the case of point 5) considering when the first choice string span was not selected, we find that only 8.5% of conversations have a string slot value that does not match the reviewer’s first choice. However, these differences are subjective and subtle, with over half of the cases concerning whether to include ‘the’ before a date (e.g. ‘the 5th of March’ vs ‘5th of March’).
@@
Appendix D Qualitative Analysis
We perform a qualitative analysis on our dev set, understanding potential limitations of our data, and suggesting ways these could be mitigated for future generation efforts.
As a related issue, the natural language response can also suggest that an intent has been performed before confirmation is given by the user. Informed by this finding, we filter out conversations when an intent was not performed, unless there was a cancellation signal provided by the user (removing 79 conversations). After this filtering, our quantitative analysis finds that 4.5% of conversations include responses that suggest an intent has been performed before confirmation is given. However, in each case there was no impact on the conversation beyond the phrasing of the natural language response. As LUCID prioritises validating the system labels, we do not implement validation checks on the natural language response. Introducing additional validation for the natural language responses is likely to also improve their quality.
Finding 2) The planner’s choice of slots and their corresponding values can sometimes be unrealistic. While a strength of LUCID is the realistic and varied slot values used in conversations, this is not always the case. We also notice that the choice of slots included in a conversation is not always realistic. For example, you would not usually give the start time, end time in addition to specifying the duration of a swimming lesson. Our quantitative analysis identifies that 6% of conversations contain at least one unrealistic slot value, while 3.5% of conversations include an unrealistic combination of slots. The unrealistic slot combinations demonstrate a limitation to our sampling approach, where we randomly sample which optional slots should be included in each conversation. This issue could be overcome with an additional LLM stage responsible for deciding if the slot combination provided is realistic or not.
Finding 4) The system command labelling is consistently high quality, with few labelling mistakes. We quantify this finding with our quantitative analysis of 200 conversations in the development set, which finds only 1% of examples when the system label is not correct. More detail on the two system labelling errors identified are provided in Appendix C.
Appendix E Validation and Post-Processing
We introduce additional validation, ensuring turns that include our challenging conversational phenomena are correctly predicted by our LLMs. When the conversation rules instruct a user to introduce a specific conversational phenomenon for a certain slot value, the user is instructed to also provide a signal (in the form of a special token) to show that this unhappy path is being performed. We then use this signal for validation purposes, ensuring that the following system command matches the expected response for this phenomenon. However, we do not provide these special tokens to the system which interprets and labels the user request; these would not be available to a real virtual assistant and we find that including them during data generation can result in unrealistic target labels (e.g. if a user’s ‘irrelevant’ response accidentally constitutes a plausible slot value).
A range of post-processing rules are also introduced after our qualitative analysis. We filter conversations where a slot is corrected during the conversation, but where there is no correctional signal provided by the user (as described above, a signal is provided by the user for each complex conversational phenomena which is used purely for validation purposes). This filtering process removes instances where a slot was first mentioned by the user without giving a value, with the system incorrectly assigning a slot value from this turn (filtering 123 conversations). We perform additional filtering to remove empty string slot values (removing 27 conversations), and any instances where the system turn predicts a hint, as hints should only occur in Signal turns. There were 172 occurrences when a hint was predicted by the system, although in almost all cases these conversations were already filtered by another post-processing filter.
In total, 56% of conversations pass all of our validation checks. To avoid wasting valuable conversational data, we salvage the prefix of an aborted conversation up to the point where the validation error was identified [^5] To avoid overly short conversations, we do this only if at least one intent has been performed already or at least 10 turns have occurred. In these cases, we truncate the conversation, sampling LLM generated natural language responses that justify interrupting the conversation.
Appendix F Annotation Schema Detail
An important part of our annotation schema is the order of the turns, and how system turns trigger the natural language responses. This section provides more detail on these points.
LUCID automatically creates the mock back-end for each intent using the schema generated in steps 1 and 2. This involves generating a Python class to represent the intent in question, which is then instantiated as an object and interacts with the system commands to indicate when mandatory slots have not yet been provided, or when confirmation is still required before the intent can be performed. The outputs of the mock back-end are represented by the signal turns described above.
Appendix G Baseline Results
@@
We train six different baseline models on the LUCID training data (T5-small, T5-base, T5-large, Flan-T5-small, Flan-T5-base and Flan-T5-large models). Each model is evaluated on the test set for seen intents and the OOD test set for unseen intents (see Table 6). We additionally experiment with training our Flan-T5-base baseline on varying amounts of training data (see Table 8). Details about about the choice of hyper-parameters can be found in Appendix K.
@@
To measure performance on our generated LUCID data, we consider five performance metrics: joint goal accuracy, intent accuracy, fuzzy slot accuracy, exact match accuracy (between user turns), and exact match accuracy for an entire dialogue.
@@
We also use fuzzy matching for our slot accuracy measure, which is a joint accuracy measure across all the slot values provided in a single system turn (when any slot values are predicted, or when they are included in the labels). We additionally introduce exact match metrics that consider the accuracy of all system commands, not just those that refer to intent and slot values (for example, including ‘say’ commands). We introduce two exact match metrics - exact match (turn) considers whether all predicted system commands between two user turns exactly match their labels, while exact match (conversation) considers whether every predicted system command in a conversation matches with the system labels.
@@
For the tool retrieval, as gold system labels for previous turns are seen in the prompt conversation history, we retrieve all tools that have been mentioned in an oracle history up to that point.
@@
Table 4: Total number of each conversational phenomenon within each split of our dataset. While there are few conversations for cancellation, this behaves similarly to the ‘delay confirmation’ phenomenon. We also show the exact match (Turn) metric for each conversational phenomena from a T5-Flan-base baseline model (details of the metric are provided in section Appendix G).
Appendix H Intent Generation
@@
A transactional intent was created for each of the remaining 53 descriptions. LUCID then creates a query intent corresponding to each transactional intent. The query intent returns entities that would be created using the corresponding transactional intent.
Some query intents were merged together (this happens when the corresponding transactional intents had the same entity names - one of the intent properties generated by the LLM). As a result, there are 6 fewer query intents than transactional intents, resulting in a total of 100 intents. Each of the following pairs of transactional intents returned the same entity names, and so their corresponding query intents were combined: add_tv_program_to_favorites and add_artist_to_favorites (both of which return ‘favorites’ entities), set_timer and set_alarm (both of which return ‘alarm’ entities), book_nails_appointment and book_spa_appointment (both of which return ‘appointments’), order_supermarket_shop and order_takeaway (both of which return ‘orders’), add_song_to_favorites and play_song (both of which return ‘songs’), and review_film and review_restaurant (both of which return ‘reviews’).
@@
Appendix I Slot Duplication within PRESTO and SGD